A computer is infected with the Sasser virus. Assumethat it infects 20 other computers within 5 minutes; and that these PCs and servers each infect 20 more machines within another 5 minutes, etc. How long until 100 million computers are infected?
Answer by stanbon(48507)

A computer is infected with the Sasser virus. Assume that it infects 20 other computers within 5 minutes; and that these PCs and servers each infect 20 more machines within another 5 minutes, etc. How long until 100 million computers are infected?
-----------------
EQUATION:
N(t) = 20*5^(x/5), where x is number of elapsed minutes.
100x10^6 = 20*5^(x/5)
5x10^6 = 5^(x/5)
Take the log of both sides to get:
log5 + 6 = (x/5)log5
x/5 = 1 + 6/log5
x = 5 + 30/log5
x = 47.92 minutes
---------------------
Cheers,
Stan H.